# BED-LLM: Intelligent Information Gathering with LLMs and Bayesian Experimental Design

### v1

@@ Abstract
We propose a general-purpose approach for improving the ability of Large Language Models (LLMs) to intelligently and adaptively gather information from a user or other external source using the framework of sequential Bayesian experimental design (BED). This enables LLMs to act as effective multi-turn conversational agents and interactively interface with external environments. Our approach, which we call BED-LLM (Bayesian Experimental Design with Large Language Models), is based on iteratively choosing questions or queries that maximize the expected information gain (EIG) about the task of interest given the responses gathered previously. We show how this EIG can be formulated in a principled way using a probabilistic model derived from the LLM’s belief distribution and provide detailed insights into key decisions in its construction. Further key to the success of BED-LLM are a number of specific innovations, such as a carefully designed estimator for the EIG, not solely relying on in-context updates for conditioning on previous responses, and a targeted strategy for proposing candidate queries. We find that BED-LLM achieves substantial gains in performance across a wide range of tests based on the 20–questions game and using the LLM to actively infer user preferences, compared to direct prompting of the LLM and other adaptive design strategies.
@@ 1 Introduction
Despite their many successes, LLMs currently fall short on a crucial aspect of interactive intelligence: proactively seeking out information from a user or external environment in an intelligent and adaptive manner (Laban et al,, 2025; Li et al, 2025c, ). For example, they have been shown to perform poorly on problems such as multi-step guessing games (Bertolazzi et al,, 2023; Zhang et al,, 2024), task clarification (Chi et al,, 2024), IT task automation (Jha et al,, 2025), and iterative external tool use (Patil et al,, 2025). In particular, while modern LLMs are often capable of producing coherent and insightful questions (or other external queries) in a one-shot manner, they typically struggle to appropriately tailor their questions to previously gathered responses on interactive tasks (Bertolazzi et al,, 2023; Patil et al,, 2025).
@@
There is, therefore, a pressing need to improve the ability of LLMs to adaptively ask questions and gather information in a targeted manner. Such capabilities are essential for a wide variety of problems, such as clarifying user intent, personalizing model behavior to a particular user, or generally acting as effective multi-turn conversational agents. They are also critical if we want to use LLMs in data gathering tasks or as automated agents in decision-making pipelines (Wu et al,, 2025). In turn, these capabilities are essential for a wide variety of downstream applications from medical diagnostic assistants, targeted code generation, and tutoring systems, to conducting automated surveys, AI–driven scientific inquiries, and market research (Liu et al, 2024a, ; Mandal et al,, 2025; Jacobsen et al,, 2025; Lu et al,, 2024). Note that in all these problems it is not enough for the LLM to simply one-shot generate good fixed questions based only on its parametric knowledge. Instead, we need it to be able to adaptively choose questions that are good, conditioned on the already-collected user responses.
@@
We propose to address this challenge using the framework of sequential Bayesian experimental design (BED, Lindley, (1956); MacKay, (1992); Chaloner & Verdinelli, (1995); Sebastiani & Wynn, (2000); Rainforth et al, (2024)), which provides a model-based, information-theoretic mechanism for making adaptive design decisions, given a generative model of the experiment. Specifically, we show how the problem of interactive information gathering with LLMs can be formulated as a sequential experimental design problem, wherein we iterate between choosing queries based on maximizing their expected information gain (EIG) and updating our beliefs with the information from the received response. Here, the underlying generative model from which the sequential BED procedure is formed is derived from the LLM, and we carefully show how this should be constructed, providing extensive insights on key design decisions.
We call our approach BED-LLM (which stands for Bayesian Experimental Design with Large Language Models) and show how its success is critically dependent on careful choices for how the EIG is estimated, how candidate queries are generated, and how the model is updated given previous responses. For example, we show how approximations to the EIG made in previous work (Kobalczyk et al,, 2025; Hu et al,, 2024) can significantly degrade performance compared to using more carefully constructed estimation strategies. We also show how simple in-context learning (Brown et al,, 2020) is typically not sufficient to incorporate all the information from previous query–response pairs, with an additional sample–then–filter strategy introduced to correct this. Finally, we introduce a targeted strategy for proposing candidate queries (whose EIG we then evaluate) that directly utilizes our current belief state and provides important improvements over direct generation by the LLM.
@@
Together, we find that these innovations provide substantial performance benefits over directly generating queries from the LLM and more basic approximations of the sequential BED framework. Specifically, we first find that BED-LLM provides substantial improvements in the success rate for the 20-Questions problem across a variety of LLMs and target quantities. For example, when predicting celebrities with Mistral-Large, we observe an improvement in success rate from 14% to 91%. Second, we demonstrate noticeable improvements in using the LLM for movie recommendations, showing that these benefits hold even when the LLM’s predictive model differs from that of the answerer.
In summary, we illustrate how the adaptive information gathering capabilities of LLMs can be significantly improved by the careful application of the sequential BED framework to their decision making, thereby making an important step towards their more general usage as interactive agents.
@@ 2 Problem Formulation and Background
There are two natural ways one might look to improve LLMs’ adaptive information gathering capabilities: we can look to change the LLM itself through finetuning or altering the pretraining process, or we can make changes in how the LLM is used at deployment time. Our focus will be on the latter, as this means we can make improvements to performance without requiring access to any additional data or expensive retraining, noting that for information–gathering problems, we will not usually have task–specific data available upfront (e.g. we cannot expect our LLM to have any knowledge about a particular user’s preferences). However, we emphasize that any improvements that can be made to the LLM itself will be complementary to our deployment–time approach.
@@
To formalize the notion of information gathering, we need a concrete idea of what we wish to learn about. We denote this target quantity as $\theta$, and note that it could, for example, be a user’s preference, the answers to one or more questions, a recommendation, or a piece of content that a user wants to generate. To meaningfully reason about $\theta$, we need to have some belief distribution for it, which we denote $p\left(\theta;h_{t}\right)$, [^1] Note we carefully distinguish between explicit conditioning, i.e. $p(a|b)$, and more general dependency, $p(a;b)$, in our notation. The former explicitly corresponds to the conditional distribution of an associated joint distribution, $p(a,b)$, while the latter may not. So here, for example, $h_{t}$ influences our distribution on $\theta$, but we are not deriving this distribution through conditioning of some explicit joint distribution. where $h_{t}$ represents the history of queries and responses at time $t$. As more queries are made, this history will grow and our beliefs on $\theta$ will update accordingly. We note that $p\left(\theta;h_{t}\right)$ will typically not be defined explicitly. Even in the LLM setting, there is significant flexibility to how it is constructed, as we discuss in Sections 3.3 and 4.
2.1 Information Gathering with LLMs
Our focus is on how to use LLMs to gather information about our target quantity $\theta$. The precise set of problems this can relate to is quite broad and includes, to name but a few, task clarification (Zhang et al, 2025a, ), preference learning (Handa et al,, 2024; Chakraborty et al,, 2024; Ouyang et al,, 2022), diagnosis and troubleshooting (Hirosawa et al,, 2024; Jha et al,, 2025), automating external simulations or real-world experiments (Mandal et al,, 2025), conducting surveys or collecting feedback (Aher et al,, 2023; Lee et al,, 2024), and AI tutoring agents (Kestin et al,, 2025; Liu et al, 2024a, ).
@@ 2.1 Information Gathering with LLMs
What all of these have in common is that the LLM starts with incomplete information about the target $\theta$, but can refine its beliefs by making queries, $x\in\mathcal{X}$, to the user or some other external agent and receiving responses, $y\in\mathcal{Y}$, that are informative about $\theta$. In general, multiple such queries, $x_{1},\dots,x_{n}$, can be made. We can choose them adaptively in a sequential decision-making process where we iteratively choose each $x_{t+1}$ based on the collected history $h_{t}:=(x_{i},y_{i})_{i=1:t}$. For clarity of exposition, we focus on the case where the $x_{i}$ correspond to explicit questions asked to the user, but emphasize that the approach applies more broadly to other forms of external interaction by the LLM, such as retrieving documents or calling external functions.
@@
The simplest and cheapest way to incorporate the interaction history into the LLM to aid our sequential decision making is to include it in the context (Brown et al,, 2020). Namely, if the LLM’s distribution over generated text, $z\in\mathcal{Z}$, is $p_{\mathrm{LLM}}(z;c)$, where $c$ is an initial prompt, then we can consider $p_{\mathrm{LLM}}(z;[c,h_{t}])$ as an updated LLM that uses the previous question–response pairs as in-context examples, where $c,h_{t}]$ represents some appropriate concatenation of the history into the context. From this, we can derive an updated belief distribution over $\theta$. Most simply, this can be done by using $p_{\mathrm{LLM}}(z;[c,h_{t}])$ to directly query about $\theta$ (e.g. if $\theta$ is some preference, we could prompt the LLM to predict this preference). However, as we show later, this approach often fails to appropriately or fully incorporate the information from $h_{t}$, leading to belief states that are inconsistent with past observations. We will introduce a more robust and sophisticated method for deriving $p(\theta;h_{t})$ from $p_{\mathrm{LLM}}(z;[c,h_{t}])$ in Section˜3.3.
In the interest of brevity, for the rest of the paper we will drop the initial prompt $c$ from our notation and also assume that an appropriate final prompt has been added to end of the context to predict/sample our quantity of interest. For example, we will use $p_{\mathrm{LLM}}(\theta;h_{t})$ to denote the distribution when our context is $c,h_{t}]$ followed by an appropriate prompt to sample $\theta$ as its next token(s).
@@ 2.2 Information–Theoretic Experimental Design
Our aim of maximizing the information gathered about a target quantity can be formalized using the framework of Bayesian experimental design (BED; Lindley,, 1956; 1972; Chaloner & Verdinelli,, 1995; Rainforth et al,, 2024), which provides a powerful and principled framework for data acquisition.
The core of the BED framework is a joint generative model, $p(\theta,y;x)$ over the target quantity $\theta$ and outcomes, $y$, given designs $x$. Traditionally, this is most commonly specified as a conventional Bayesian model using a prior $p\left(\theta\right)$ and likelihood $p(y|\theta;x)$ (Rainforth et al,, 2024), but it can also be directly defined using a stochastic simulator (Kleinegesse & Gutmann,, 2020; Rainforth,, 2017) or as the marginal of some larger model (e.g. because $\theta$ corresponds to something other than the model parameters (Bickford Smith et al,, 2023; MacKay,, 1992)).
@@
In the most general case, designs are then chosen to maximize the expectation of some utility function $U(\theta,y,x)$ under this generative model; that is we take $x^{*}=\mathrm{argmax}_{x}\mathbb{E}_{p(\theta,y;x)}\left[U(\theta,y,x)\right]$. The most common choice is to take $U(\theta,y,x)=\log p(\theta,y;x)-\log p\left(\theta\right)\log p(y;x)$ where $p\left(\theta\right)$ and $p(y;x)$ are the marginal distributions on $\theta$ and $y$ implied by our joint model and we have assumed that our current beliefs on $\theta$ are independent of the design. This leads to an objective corresponding to the expected information gain (EIG) in $\theta$ (Lindley,, 1956; 1972),
$\mathrm{EIG}_{\theta}(x)$ $=\mathbb{E}_{p(\theta,y;x)}\left[\log\frac{p(\theta,y;x)}{p\left(\theta\right)p(y;x)}\right],$ (1)
which using Bayes’ rule can also be represented in the following equivalent forms
$\text{EIG}_{\theta}(x)$ $=\mathrm{H}[p\left(\theta\right)]-\mathbb{E}_{p(y;x)}[\mathrm{H}[p(\theta|y;x)]]$ (2)
$=\mathrm{H}[p(y;x)]-\mathbb{E}_{p\left(\theta\right)}[\mathrm{H}[p(y|\theta;x)]],$ (3)
@@
where $\mathrm{H}$ denotes the Shannon entropy (i.e., $\mathrm{H}[p\left(\theta\right)]=-\mathbb{E}_{p\left(\theta\right)}[\log p\left(\theta\right)]$). We can thus equivalently think of this EIG as the mutual information between $\theta$ and $y$, the expected information gain over possible data simulated from our model (where the information gain is defined as the reduction in entropy from our prior on $\theta$ to the posterior), and the expected reduction in entropy over data from observing $\theta$ simulated from our prior (Sebastiani & Wynn,, 2000). By maximizing this objective with respect to $x$, we are thus maximizing the expected amount of information we gain in $\theta$, subject to the assumptions that our generative model can both accurately estimate the true data-generating process and encapsulate our beliefs about $\theta$.
Working with the EIG is highly suited to a sequential or adaptive design approach, wherein it is generally referred to as sequential BED or Bayesian Adaptive Design (Rainforth et al,, 2024). Because the EIG is only a function of our underlying model, when we update the model as new data becomes available, our EIG design objective will naturally update as well. In particular, if $p(\theta,y_{t+1};h_{t},x_{t+1})$ represents our generative model after already observing the history $h_{t}$, then we can iteratively optimize the incremental EIG (MacKay,, 1992; Cavagnaro et al,, 2010),
$\mathrm{EIG}_{\theta}(x_{t+1};h_{t})$ $=\mathbb{E}_{p(\theta,y_{t+1};h_{t},x_{t+1})}\left[\log\frac{p(\theta,y_{t+1};h_{t},x_{t+1})}{p\left(\theta;h_{t}\right)p(y_{t+1};h_{t},x_{t+1})}\right].$ (4)
Here our updated model, $p(\theta,y_{t+1};h_{t},x_{t+1})$, conventionally comes from a Bayesian update of the original model, noting that the Savage axioms dictate that a rational agent should update its beliefs in a Bayesian manner (Savage,, 1954). However, in many cases, this is not practical and other non-Bayesian updates are performed instead. In particular, in many Bayesian active learning problems, the update actually corresponds to retraining the model with an updated dataset (Gal et al,, 2017; Bickford Smith et al,, 2023).
It has recently been noted that this traditional sequential BED approach can sometimes be suboptimal in practice, as it only optimizes the EIG of the next observation, without planning ahead for the fact that design decisions taken at a given step can also influence the achievable EIGs from future steps (Foster,, 2021). A variety of policy–based BED approaches have subsequently been proposed to address this (Foster et al,, 2021; Ivanova et al,, 2021; Blau et al,, 2022; Huan & Marzouk,, 2016; Hedman et al,, 2025), while also removing the need to make model updates and conduct optimizations during the experiment itself. However, training these policies typically requires a very large number of simulations from the generative model, which makes them impractical for our setting, due to the expense of simulating full experimental rollouts from the LLM.
@@ 3 Sequential Bayesian Experimental Design for LLMs
The sequential BED framework described in Section˜2.2 requires two core components to be specified by the user: an initial joint model $p(\theta,y;x)$ over hypotheses $\theta$ and outcomes $y$, given chosen experiment(s) $x$; and a procedure to derive an updated model $p(\theta,y_{t+1};h_{t},x_{t+1})$ after observing an experimental history $h_{t}$. In the LLM setting, there is significant flexibility in these critical design decisions. In particular, there are various ways that one can derive a suitable joint distribution from the LLM and the ability to learn in context provides opportunities for update methods that go beyond standard Bayesian model updates.
@@
A major challenge in the LLM setting is that unlike conventional probabilistic models, in general, $p_{\mathrm{LLM}}(\theta)\,p_{\mathrm{LLM}}(y;[\theta,x])\neq p_{\mathrm{LLM}}(y;x)\,p_{\mathrm{LLM}}(\theta;[x,y])$. That is, we induce a different joint distribution if we first sample $\theta$ then sample $y$ with $\theta$ in context (which we refer to as the prior–likelihood pairing), than if we first sample $y$ then sample $\theta$ with $y$ in context (which we refer to as the data–estimation pairing). Moreover, in each pairing we may wish to deviate from the LLM’s directly induced distribution on one or both components. The success of using sequential BED with LLMs turns out to be critically dependent on which pairing we choose and the deviations we make from LLM’s directly induced distribution.
We delay proper discussion of this complex issue until Section˜4, where we will see that the preferable setup can depend on the precise problem setting and, in particular, the relative complexity of spaces of $\theta$ and $y$. For now, we will focus on using the prior–likelihood pairing, as this transpires to generally be the advantageous setup when the space of $\theta$ is more complex than that of $y$, which we believe to be the more commonly occurring setting and the one we consider in our experiments. While we will generally use the LLM’s directly induced distribution for the likelihood, we allow the prior to deviate from this in a problem–specific manner. As such, our initial joint model will be $p(\theta,y;x)=p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])$, where the specification of the prior $p(\theta)$ can vary between problems.
@@
Optimally updating the joint model in this setting requires incorporating new observations in a way that both fully captures the available information from new data and is computationally tractable. At one extreme, we could perform full Bayesian updates via approximate inference, as in classical sequential BED. This approach, however, demands a prohibitively large number of LLM evaluations to accurately approximate the posterior, and it does not exploit the power of the LLM as a probabilistic generative model, where autoregressive sequential rollouts often lead to more nuanced and diverse behavior than repeated static likelihood queries. The other obvious choice is to rely on simple in context updates, such that we use $p(\theta;h_{t})=p_{\mathrm{LLM}}(\theta;h_{t})$. However, as we show later, this fails to sufficiently capture the information provided by the new data and thus undermines the sequential BED approach.
As we will discuss in Section˜3.3, we therefore employ a strategy that is somewhere between the two, wherein we set up a rejection sampling procedure that draws samples from the in–context distribution, $p_{\mathrm{LLM}}(\theta;h_{t})$, but then filters out samples that are actually not compatible with previous observations. We refer to the resulting distribution as $p_{f}(\theta;h_{t})$.
In our experiments and subsequent exposition of our approach, we will not update the likelihood model as more data is gathered, that is, our likelihood in the sequential setting will be $p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ instead of $p_{\mathrm{LLM}}(y_{t+1};[h_{t},\theta,x_{t+1}])$. The main rationale of this choice is that for many problems our beliefs on $\theta$ capture all the required information to predict $y|x$, hence including the history is adding unnecessary context that could influence the LLM’s behavior in undesirable ways. However, it is important to note that $p_{\mathrm{LLM}}(y_{t+1};[h_{t},\theta,x_{t+1}])$ should be used instead for problems where $\theta$ will not capture all information from previous data, e.g. if $\theta$ is a binary value corresponding to whether we reject a null hypothesis, or is the answer to a particular other question of interest.
@@
We now introduce our specific principled algorithm for intelligent and adaptive information gathering in LLMs using sequential BED. We call our approach BED-LLM, which stands for Bayesian Experimental Design with Large Language Models. Here, the questions posed to the user [^2] More broadly, these could be more general queries such as document retrieval or function calling; however in our experiments we only consider questions asked of a user. will correspond to our designs, $x$, and the responses received will correspond to our outcomes, $y$. Using the LLM to derive a joint model over these outcomes and the target variables $\theta$ as described above, we can derive a precise form of the expected information gain, $\text{EIG}_{\theta}(x)$. We can then interleave choosing informative questions by optimising the incremental EIG defined in Equation˜4, and updating our underlying model based on the received question–response pairs. Specifically, BED-LLM iterates over the following key steps, where $t$ indexes the current turn.
@@
    Generate Candidate Questions: Propose a candidate set of $M$ diverse, multiple-choice questions, $\mathcal{X}^{\text{cand}}$, by appropriate sampling of the LLM based on the current conversational context (Section˜3.1).
@@
    Compute EIG Estimator: For each candidate question $x\in\mathcal{X}^{\text{cand}}$, estimate $\text{EIG}_{\theta}(x)$ as described in in Section˜3.2.

### v2

@@ Abstract
We propose a general-purpose approach for improving the ability of large language models (LLMs) to intelligently and adaptively gather information from a user or other external source using the framework of sequential Bayesian experimental design (BED). This enables LLMs to act as effective multi-turn conversational agents and interactively interface with external environments. Our approach, which we call BED-LLM (Bayesian experimental design with large language models), is based on iteratively choosing questions or queries that maximize the expected information gain (EIG) about the task of interest given the responses gathered previously. We show how this EIG can be formulated (and then estimated) in a principled way using a probabilistic model derived from the LLM’s predictive distributions and provide detailed insights into key decisions in its construction and updating procedure. We find that BED-LLM achieves substantial gains in performance across a wide range of tests based on the 20 Questions game and using the LLM to actively infer user preferences, compared to direct prompting of the LLM and other adaptive design strategies.
@@ 1 Introduction
Intelligent information gathering—the ability to ask the right questions at the right time—is fundamental to effective AI systems. However, despite their many successes, LLMs currently fall short on a crucial aspect of interactive intelligence: proactively seeking out information from a user or external environment in an intelligent and adaptive manner (Laban et al., 2025; Li et al., 2025c). For example, they have been shown to perform poorly on problems such as multi-turn guessing games (Bertolazzi et al., 2023; Zhang et al., 2024), task clarification (Chi et al., 2024), IT task automation (Jha et al., 2025), and multi-step tool use (Patil et al., 2025). In particular, while modern LLMs are often capable of producing coherent and insightful questions (or other external queries) in a single-turn setting, they typically struggle to appropriately tailor their questions to previously gathered responses on interactive tasks (Bertolazzi et al., 2023; Patil et al., 2025).
@@
There is, therefore, a pressing need to improve the ability of LLMs to adaptively ask questions based on previous responses, and gather information in a targeted manner. Such capabilities are essential for a wide variety of problems, such as clarifying user intent, personalizing model behavior to a particular user, or generally acting as effective multi-turn conversational agents. They are also critical if we want to use LLMs in data gathering tasks or as automated agents in decision-making pipelines (Wu et al., 2025). In turn, these capabilities are essential across domains ranging from medical diagnosis (Hirosawa et al., 2024), troubleshooting (Jha et al., 2025), preference learning (Handa et al., 2024; Chakraborty et al., 2024; Ouyang et al., 2022), and tutoring systems (Kestin et al., 2025; Liu et al., 2024a), to conducting automated surveys (Aher et al., 2023; Lee et al., 2024; Jacobsen et al., 2025), and AI–driven scientific inquiries (Lu et al., 2024; Mandal et al., 2025). Note that in all these problems it is not enough for the LLM to generate full sets of suitable questions up front, we need it to be able to adaptively choose questions that are tailored to the already-collected user responses.
@@
We propose to address this challenge using the framework of sequential Bayesian experimental design (BED, Lindley (1956); MacKay (1992); Chaloner & Verdinelli (1995); Sebastiani & Wynn (2000); Rainforth et al. (2024)), which provides a model-based, information-theoretic mechanism for making adaptive design decisions, given a generative model of the experiment. Specifically, we show how the problem of interactive information gathering with LLMs can be formulated as a sequential experimental design problem with a model derived from the LLM, wherein we iterate between choosing queries based on maximizing their expected information gain (EIG) and updating our beliefs with the information from the received response.
We call our approach BED-LLM and show how its success is critically dependent on our precise model formulation, belief updating procedure, and EIG estimation strategy. In particular, we show that it is essential to formulate the model with a precise distribution pairing that does not solely rely on in-context learning to update beliefs and uses the LLM’s uncertainties in the space of answers rather than the more complicated underlying hypothesis space we are trying to learn in.
@@
Together, we find that these innovations provide substantial performance benefits over directly generating queries from the LLM and more basic approximations of the sequential BED framework. Specifically, we first find that BED-LLM provides substantial improvements in the success rate for the 20 Questions problem across a variety of LLMs and target quantities. For example, when guessing celebrities with a small model, Llama-3.1-8B, we observe a 5.8x gain in success rate. Second, we demonstrate noticeable improvements in using the LLM for movie recommendations, showing that these benefits hold even when the LLM’s predictive model differs from that of the answerer.
@@ 2 Problem Formulation and Background
There are two natural ways to improve LLMs’ ability to gather information: modifying the model itself (e.g. via test-time- or post-training) or altering how the model is used at deployment time. We focus on the latter, since information–gathering tasks rarely provide task-specific data upfront (e.g. a user’s unknown preferences), and deployment–time methods avoid the cost and difficulty of finetuning an LLM altogether and are applicable to any existing LLM. However, we emphasize that improvements at the model level (e.g., Zhang et al., 2024) would be complementary to our approach.
@@
To formalize the notion of information gathering, we need a concrete idea of what we wish to learn about. We denote the target quantity of interest as $\theta$, which may represent, for example, a user’s preferences, the answer to a question, or a desired piece of content. We start with incomplete information about $\theta$, as represented by an initial belief distribution or prior, $p(\theta)$, but can refine these beliefs by making queries, $x\in\mathcal{X}$, to the user or some other external agent and receiving responses, $y\in\mathcal{Y}$, that are informative about $\theta$. Multiple such queries, $x_{1},\dots,x_{n}$, can be adaptively selected in a sequential decision-making process where we iteratively choose each $x_{t}$ based on the collected history $h_{t-1}:=(x_{i},y_{i})_{i=1:t-1}$. As our history grows, we will update our belief distribution to obtain $p(\theta;h_{t-1})$ via some model updating procedure. [^1] We carefully distinguish between explicit probabilistic conditioning, i.e. $p(a|b)$, and more general dependency, $p(a;b)$. The former corresponds to the conditional distribution of an associated joint distribution, $p(a,b)$, while the latter may not. Here, $h_{t-1}$ influences our distribution on $\theta$, but it is not derived via a joint distribution. In the LLM setting, there is considerable flexibility in how $p(\theta;h_{t-1})$ is constructed, as discussed in §3.3 and §4. While $p(\theta;h_{t-1})$ need not be explicitly defined, it provides the foundation for our information-theoretic method of query selection.
@@
For clarity of exposition, we focus on the case where the $x_{t}$ correspond to explicit questions asked to the user, but emphasize that the approach applies more broadly to other forms of external interaction by the LLM, such as retrieving documents or calling external functions.
2.1 In–Context Updating of the belief distribution
@@
A natural and cheap way to incorporate the interaction history into the LLM is to include it in the context (Brown et al., 2020). If the LLM’s distribution over generated text, $z\!\in\!\mathcal{Z}$, is $p_{\mathrm{LLM}}(z)$ given appropriate prompting, then $p_{\mathrm{LLM}}(z;h_{t-1})$ is an updated distribution with the previous question–response pairs in context. From this, we can derive an updated belief distribution over $\theta$. Most simply, this can be done by using $p_{\mathrm{LLM}}(z;h_{t-1})$ to directly query about $\theta$ (e.g. if $\theta$ is some preference, we could prompt the LLM to predict this preference). However, as we show later, this approach often fails to appropriately incorporate the information from $h_{t-1}$, leading to a belief distribution inconsistent with past observations. This is consistent with recent work that shows that in context updating does not treat all contextual information equally (Kossen et al., 2024; Liu et al., 2024b; Zhang et al., 2024). In §3.3, we introduce a more robust method for deriving $p(\theta;h_{t-1})$.
@@ 2.2 Information–Theoretic Experimental Design
The core of the BED framework is a joint generative model $p(\theta,y;x)$ over the target quantity $\theta$ and outcomes, $y$, given designs $x$. Most commonly, this is specified as a Bayesian model using a prior $p(\theta)$ and likelihood $p(y|\theta;x)$. In the general case, designs are then chosen to maximize the expectation of some utility function $U(\theta,y,x)$ under this model: we choose $x^{*}=\mathrm{argmax}_{x}\mathbb{E}_{p(\theta,y;x)}\left[U(\theta,y,x)\right]$. The most common choice is to take $U(\theta,y,x)=\log p(\theta,y;x)-\log p(\theta)\log p(y;x)$, where $p(\theta)$ and $p(y;x)$ are the marginal distributions on $\theta$ and $y$ implied by our joint model and we have assumed that our current beliefs on $\theta$ are independent of the design $x$. This leads to an objective corresponding to the expected information gain (EIG) in $\theta$ (Lindley, 1956; 1972),
$\text{EIG}_{\theta}(x)$ $=\mathrm{H}[p(\theta)]-\mathbb{E}_{p(y;x)}[\mathrm{H}[p(\theta|y;x)]]$ (1)
$=\mathrm{H}[p(y;x)]-\mathbb{E}_{p(\theta)}[\mathrm{H}[p(y|\theta;x)]],$ (2)
@@
where $\mathrm{H}$ denotes the Shannon entropy (i.e., $\mathrm{H}[p(\theta)]=-\mathbb{E}_{p(\theta)}[\log p(\theta)]$). We can thus equivalently think of the EIG as: a) the mutual information between $\theta$ and $y$, b) the expected information gain over possible data simulated from our model (where the information gain is defined as the reduction in entropy from our prior on $\theta$ to the posterior), or c) the expected reduction in entropy over data from observing $\theta$ simulated from our prior (Sebastiani & Wynn, 2000).
Working with the EIG is highly suited to a sequential or adaptive design approach, wherein it is generally referred to as sequential BED or Bayesian Adaptive Design (Rainforth et al., 2024). Because the EIG is only a function of our underlying model, when we update the model as new data becomes available, our EIG design objective will naturally update as well. Specifically, to derive the incremental EIG (Cavagnaro et al., 2010) for the $t$-th query, $\text{EIG}_{\theta}(x_{t};h_{t-1})$, we simply replace the joint $p(\theta,y;x)$ in the above formulation with the updated joint $p(\theta,y_{t};h_{t-1},x_{t})$, with all marginals a conditionals derived from this (e.g. $p(y;x)$ becomes $p(y_{t};h_{t-1},x_{t})$). Here this updated joint conventionally comes from a Bayesian update of the original model. However, in many cases, this is not practical and other non-Bayesian updates are performed instead, e.g. in active learning the update often actually corresponds to retraining the model with the new data (Gal et al., 2017; Bickford Smith et al., 2023).
@@ 3 Sequential Bayesian Experimental Design for LLMs
The sequential BED framework described in §2.2 requires two core components to be specified by the user: a) an initial joint model $p(\theta,y;x)$ over hypotheses $\theta$ and outcomes $y$, given chosen experiment $x$, and b) a procedure to derive an updated model $p(\theta,y_{t};h_{t-1},x_{t})$ after observing $h_{t-1}$. In the LLM setting, there is significant flexibility in these critical design decisions. In particular, there are many ways to derive a suitable joint distribution from the LLM and its ability to learn in-context provides opportunities for update methods that go beyond standard Bayesian model updates.
@@
A major challenge in the LLM setting is that unlike conventional probabilistic models, in general, $p_{\mathrm{LLM}}(\theta)\,p_{\mathrm{LLM}}(y;[\theta,x])\neq p_{\mathrm{LLM}}(y;x)\,p_{\mathrm{LLM}}(\theta;[x,y])$. That is, we induce a different joint distribution if we first sample $\theta$ then sample $y$ with $\theta$ in context (which we refer to as the prior–likelihood pairing), than if we first sample $y$ then sample $\theta$ with $y$ in context (data–estimation pairing). Moreover, we can deviate from the distribution directly induced by the LLM on one or both variables. The success of using BED with LLMs turns out to be critically dependent on these choices.
We delay proper discussion of this complex issue until §4, where we will see that the preferable setup can depend on problem setting and, in particular, the relative complexity of spaces of $\theta$ and $y$. For now, we will focus on using the prior–likelihood pairing; we will argue in §4 that this is the advantageous setup in many practical scenarios. While we will generally use the LLM’s directly induced distribution for the likelihood, we allow the prior to deviate from this in a problem–specific manner. As such, our initial joint model will be $p(\theta,y;x)=p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])$.
@@
Optimally updating the joint model in this setting requires incorporating new observations in a way that both fully captures the information from new data and is computationally tractable. At one extreme, we could perform full Bayesian updates via approximate inference, as in classical sequential BED. However, this demands a prohibitively large number of LLM evaluations to accurately approximate the posterior, and it does not exploit the power of the LLM as a probabilistic generative model, where autoregressive sequential rollouts often lead to more nuanced and diverse behavior than repeated static likelihood queries. At the other extreme, simple in-context updating, $p(\theta;h_{t-1})=p_{\mathrm{LLM}}(\theta;h_{t-1})$, is cheap but, as we show later, fails to reliably capture information from new data, leading to inconsistent belief states and undermining the sequential BED approach. As we discuss in §3.3, we therefore employ a strategy that is somewhere between the two: drawing samples in a way that utilizes $p_{\mathrm{LLM}}(\theta;h_{t-1})$ while encouraging diversity, then filtering out samples that are actually not compatible with $h_{t-1}$ and renormalizing. We refer to the resulting distribution as $p_{f}(\theta;h_{t-1})$. We do not update our likelihood model $p_{\mathrm{LLM}}(\theta;h_{t-1})$; see Section˜A.1 for a discussion.
@@
We now introduce our specific algorithmic approach, BED-LLM. Here, the queries will correspond to our designs, $x$ (assumed to be in form of questions posed to the user in the following for simplicity, but could also be, e.g., external function calls, document retrieval, web search, etc), and the responses received will correspond to our outcomes, $y$. Using the LLM to derive joint models over these outcomes and the target variables $\theta$ given histories, $h_{t-1}$ as described above, we can interleave choosing informative questions by optimizing the incremental EIG, $\text{EIG}_{\theta}(x_{t};h_{t-1})$, and updating our underlying model based on the received question–response pairs. Specifically, BED-LLM iterates over the following key steps, where $t$ indexes the current turn:
@@
    Generate candidate questions3.1): Propose a candidate set of $M$ diverse, multiple-choice questions, $\mathcal{X}^{\text{cand}}$, by appropriate sampling of the LLM based on the conversational context $h_{t-1}$.
@@
    Compute EIG estimator3.2): For each candidate $x_{t}\in\mathcal{X}^{\text{cand}}$, estimate $\text{EIG}_{\theta}(x_{t};h_{t-1})$.